After taking into consideration a couple's AGI in 2019, the couple's American opportunity credit and lifetime learning credit are $5,000 and $1,800, respectively. The couple's refundable and nonrefundable education credit is:

1. $5,000 and $1,800

2. $0 and $6,800

3. $3,000 and $3,800

4. $6,800 and $0

5. $2,000 and $4,800
Answer #1

5. $2,000 and $4,800 respectively.

Explanation: American opportunity tax credit is allowed maximum of $2,500 per student and maximum of 40% is refundable. Lifetime learning credit is non refundable.

Amount of AOTC = $2500*2 = $5,000

Refundable = 5000 x 40% = 2000

Non refundable = 3000+1800 = 4800